1. Consider a hypothetical economy in which there are two types of good, guns and butter. There are 4 workers of the economy. A worker can produce in either the guns industry or the butter industry.

Following table shows how much of each good can be produced per week. GUNS BUTTER

   Employment

Output Employment Output

 4 35 0 0 3 32 1 19 2 27 2 27 1 20 3 34 0 0 4 40

a) Draw society’s production possibility frontier (PPF). (2)

b) Indicate an inefficient point. Why that point inside the PPF is inefficient? (1)

c) Indicate an efficient point. Why that point on the PPF is efficient? (1)

d) Indicate an unattainable point. Explain why that point outside the frontier is unattainable.

(1)

 Page 1 of 2


2. a) What is price elasticity of demand (Ep)? Draw the demand curve, when the price elasticity is zero (Ep = 0) and infinity (Ep = ∞). What do these demand curves imply? (2)

b) When the price of a particular flavor of Pizza is tk. 400, buyer want to purchase 8000 units per year, but when the price falls to tk. 260 the quantity demanded rises to 10000 Pizza’s per year. Calculate the price elasticity of demand for Pizza and make comments your result.(3)

3. The following is the hypothetical supply and demand schedule for sugar:

  Price (Tk.)

4

8 12 16 20

Quantity demanded (Kg)

1000

800

600

400

200

Quantity Supplied (Kg)

400

400

400

400

400

         .

a) b) c)

Plot the demand and supply curve for sugar

Find the equilibrium price and quantity of sugar (1)

What will happen to equilibrium price and quantity if incomes of the consumers of sugar are increased?. Discuss and show the change graphically.(2)


1
Expert's answer
2021-03-04T15:00:50-0500


2.(a) Price elasticity of demand is a measure of the responsiveness of quantity demanded to a change in price.

When the price elasticity is zero, the graph is a vertical line as shown below.





When the price elasticity coefficient is infinity, the graph is an horizontal line as shown below





A price elasticity coefficient equal to zero indicates perfectly inelastic demand. This means that demand for a good does not change in response to price.

When the Price elasticity coefficient is equal to infinity, demand is said to be perfectly elastic. When demand is perfectly elastic, buyers will only buy at one price and no other.

(b) %change in quantity demanded=(8000-10000)*100/8000=-25
% change in price =(400-260)*100/400=35
Price elasticity=-25/35=-0.7143
This price Elasticity coefficient is less than one and is the reason the change in quantity demanded is proportionately small(8000-10000)